OpenStudy (anonymous):

okay so i have a problem please check to see if im right. i have two points. point A (-1, -1) and point B (1, 5). i had to find the slop and i got 3... is that right?????

OpenStudy (anonymous):

I'LL BE Y'ALL BEST FRIEND FOREVEER

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (-1,-1) x1 = -1 and y1 = -1 (1,5) x2 = 1 and y2 = 5 now we sub slope(m) = (5 - (-1) / (1 - (-1) slope(m) = (5 + 1) / (1 + 1) slope(m) = 6/2 = 3 yep...you got it correct

OpenStudy (anonymous):

thank you again

OpenStudy (texaschic101):

happy to help :)